OpenStudy (anonymous):

Evaluate the function f (x) = 4 • 7x for x =-1 and x = 2.

OpenStudy (taramgrant0543664):

f(x)=4*7(-1)=? f(x)=4*7(2)=? You can plug it into your calculator to get your answer

OpenStudy (anonymous):

@taramgrant0543664 is correct. Just plug in the numbers.

OpenStudy (anonymous):

would the answers just be numbers?

OpenStudy (taramgrant0543664):

Ya all you need to do is put your values of x into the x spot, each time, of your equation to solve.

OpenStudy (anonymous):

The answer would be like this I will do the one where x = -1 f(x)=4*7(-1)= -28 So f(x) = -28 when x = -1 OR f(-1) = -28 There are a few ways to write it.

OpenStudy (anonymous):

Thank you!